Circular No. 8830

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2007 H1 (McNAUGHT) R. H. McNaught reports his discovery of a new comet on CCD images

taken with the 0.5-m Uppsala Schmidt telescope in the course of the Siding Spring Survey (discovery observation tabulated below), the object then reported as having a 12′′ coma and a 25′′ tail in p.a. 250^{o}. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, several CCD astrometric observers have commented on the cometary appearance of this object. G. Hug (Scranton, KS, 0.30-m Schmidt-Cassegrain reflector) reports that his images from Apr. 18.44 UT show the comet to be slightly diffuse. E. Guido and G. Sostero, remotely using a 0.25-m f/3.4 reflector at Mayhill, NM, write that the co-adding of twenty 60-s unfiltered exposures, taken on Apr. 18.46–18.47, reveals the presence of a coma nearly 20′′ in diameter and a broad tail nearly 1′ long toward p.a. 250^{o}; co-adding ten 60-s unfiltered exposures taken on Apr. 19.47 shows a coma that is nearly 17′′ in diameter (total mag ∼ 16.3) and a tail almost 45′′ long toward p.a. 255^{o}. N. Teamo and S. Hönig report that images obtained with a 0.41-m f/8 reflector at Punaauia, Tahiti, on Apr. 19.6 show a 14′′ coma and a 2′.2 tail in p.a. 252^{o}.

2007 UT  α_{2000}  δ_{2000}  Mag.
Apr. 17.78663  21^{h}59^{m}34.^{s}74  −9^{o}09′15′′.9  16.1

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2007-H32.

T = 2007 June 15.828 TT  ω = 181.^{o}324
Ω = 147.663 }2000.0
q = 2.61460 AU  i = 11.002

COMET C/2007 E2 (LOVEJOY) Further visual total-magnitude estimates (cf. IAUC 8820): Mar. 24.66

UT, 8.7 (J. Drummond, Gisborne, New Zealand, 8-cm refractor): Apr. 12.17, 7.7 (J. J. Gonzalez, Asturias, Spain, 25×100 binoculars); 16.67, 7.9 (S. Rae, Hamilton, New Zealand, 10×50 binoculars); 18.68, 8.0 (Rae).

COMET 2P/ENCKE Visual total-magnitude estimates with 25×100 binoculars: Apr. 13.85

UT, 7.7 (J. J. Gonzalez, Asturias, Spain); 14.50, 7.4 (M. Lehký, Bacetin, Czech Republic); 17.85, 6.7 (Gonzalez).

2007 April 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green